# Bridging Fairness and Explainability: Can Input-Based Explanations Promote Fairness in Hate Speech Detection?

### v1

@@ Abstract
Natural language processing (NLP) models often replicate or amplify social bias from training data, raising concerns about fairness. At the same time, their black-box nature makes it difficult for users to recognize biased predictions and for developers to effectively mitigate them. While some studies suggest that input-based explanations can help detect and mitigate bias, others question their reliability in ensuring fairness. Existing research on explainability in fair NLP has been predominantly qualitative, with limited large-scale quantitative analysis. In this work, we conduct the first systematic study of the relationship between explainability and fairness in hate speech detection, focusing on both encoder- and decoder-only models. We examine three key dimensions: (1) identifying biased predictions, (2) selecting fair models, and (3) mitigating bias during model training. Our findings show that input-based explanations can effectively detect biased predictions and serve as useful supervision for reducing bias during training, but they are unreliable for selecting fair models among candidates.
@@ 1 Introduction
Meanwhile, current NLP models are mostly based on black-box neural networks. Despite their strong capacities, the complex architecture and large number of parameters of these models make it hard for humans to understand their behaviors (Bommasani et al., 2021). To understand neural NLP models, different types of explanations have been devised, such as input-based explanations (Yin and Neubig, 2022; Deiseroth et al., 2023; Madsen et al., 2024; Wang et al., 2025a), natural language explanations (Ramnath et al., 2024; Wang et al., 2025b), and concept-based explanations (Yu et al., 2024; Raman et al., 2024). Among these, input-based explanations, often referred to as rationales, indicate the contribution of each token to models’ predictions, and thus provide the most direct insights into models’ behaviors (Arras et al., 2019; Atanasova et al., 2022; Lyu et al., 2024).
@@
Our results show that input-based explanations can effectively detect biased predictions and help reduce bias during model training, but they are less reliable for automatic fair model selection. Furthermore, our analyses indicate that explanation-based bias detection remains robust even when models are trained to reduce reliance on sensitive features, and that these explanations outperform LLM judgments in identifying bias.
@@ 2 Related Work
Input-based explanations in NLP models aim to attribute model predictions to each input token (Lyu et al., 2024). They can be broadly categorized based on how they generate explanations: gradient-based (Simonyan et al., 2014; Kindermans et al., 2016; Sundararajan et al., 2017; Enguehard, 2023), propagation-based (Bach et al., 2015; Shrikumar et al., 2017; Ferrando et al., 2022; Modarressi et al., 2022, 2023), perturbation-based (Li et al., 2016; Ribeiro et al., 2016; Lundberg and Lee, 2017; Deiseroth et al., 2023), and attention-based methods (Bahdanau et al., 2015; Abnar and Zuidema, 2020). While most prior work has focused on encoder-only models, recent studies have also explored explaining the behaviors of generative models (Yin and Neubig, 2022; Ferrando et al., 2022; Enouen et al., 2024; Cohen-Wang et al., 2024).
@@
Another line of research focuses on mitigating bias with explanations (Dimanov et al., 2020; Kennedy et al., 2020; Rao et al., 2023). For example, Hickey et al. (2020) improves fairness by reducing reliance on sensitive features during training with SHAP explanations. Bhargava et al. (2020) and González-Silot et al. (2025) first identify predictive sensitive features using LIME and SHAP, respectively, and then remove them prior to model training. In a related approach, Grabowicz et al. (2022) traces unfairness metrics back to input features and adjusts them to mitigate bias.
However, recent research has challenged the assumption that input-based explanations can be reliably used to detect and mitigate bias. First, current explanation methods may be unfaithful, meaning that they may not always reflect the true decision-making process of models (Kindermans et al., 2016; Jain and Wallace, 2019; Ye et al., 2025). This makes it difficult to reliably detect the use of sensitive features in predictions. Second, efforts to reduce the influence of sensitive features can lead to unintended consequences, sometimes degrading both model performance and fairness (Dimanov et al., 2020). Finally, models can be deliberately trained to assign lower importance to sensitive features, thereby masking biased predictions when explanations are inspected (Dimanov et al., 2020; Slack et al., 2020; Pruthi et al., 2020).
Despite growing interest in this topic, most existing work remains qualitative or restricted to limited setups (Balkir et al., 2022; Deck et al., 2024). To the best of our knowledge, this is the first study to quantitatively and comprehensively examine the relationship between explainability and fairness in NLP models. We focus on hate speech detection as a particularly critical application. Prior research has shown that biased NLP models often rely on demographic information such as race and gender, leading to inferior performance on marginalized groups in this task (Sap et al., 2019; Mathew et al., 2021). Detecting and mitigating such biased behaviors are therefore essential to ensuring equitable opportunities for all social groups to voice their perspectives on social media.
@@ 3 Experimental Setup
An input-based explanation assigns an attribution score to each token in $\mathbf{x}$ for class $c$: $a_{1}^{c},a_{2}^{c},\dots,a_{n}^{c}$}, indicating their contribution to the prediction of class $c$. The attribution scores on the sensitive tokens, $a_{g_{1}}^{c},a_{g_{2}}^{c},\dots,a_{g_{m}}^{c}$}, are referred to as sensitive token reliance scores. If multiple sensitive tokens occur in a sentence, we take the score with the maximum absolute value as the reliance score for that example: [^1] We have also experimented with normalizing feature importance scores but found that using raw scores yielded the best results.
@@
We use two hate speech detection datasets: Civil Comments (Borkan et al., 2019) and Jigsaw (cjadams et al., 2019). To ensure coverage, we focus on three bias types and their associated groups: race (black / white), gender (female / male), and religion (Christian / Muslim / Jewish). We include examples containing identity-marking terms but exclude those with derogatory or slur-based references, as the latter can reasonably serve as direct evidence for toxic predictions. The sensitive token vocabulary is derived from Caliskan et al. (2017) and Wang and Demberg (2024). Further details on dataset pre-processing are provided in Appendix A.
@@
We evaluate two types of commonly used NLP models: encoder-only models (BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)) and decoder-only large language models (Llama3.2-3B-Instruct (Dubey et al., 2024) and Qwen3-4B (Yang et al., 2025)). We fine-tune encoder-only models on data subsets that either target a single bias type or combine all bias types. For decoder-only models, we use an instruction-based setup where the model is prompted to decide whether a test example contains hate speech. The prompt includes the definition of hate speech, the test example, and a corresponding question. As a baseline, we adopt the zero-shot setting as the default configuration.
Beyond conventional fine-tuning and prompting, we also investigate the interaction between explainability and fairness in debiased models. For encoder-only models, we apply pre-processing techniques such as group balance (Kamiran and Calders, 2012), group-class balance (Dixon et al., 2018), and counterfactual data augmentation (CDA, Zmigrod et al., 2019), as well as in-processing techniques including dropout (Webster et al., 2020), attention entropy (Attanasio et al., 2022), and causal debias (Zhou et al., 2023). For decoder-only models, we incorporate bias reduction through prompt design, including few-shot, fairness imagination (Chen et al., 2025), and fairness instruction prompting (Chen et al., 2025). Further details are provided in Appendix A.
@@
For both types of metrics, higher scores indicate more bias in model predictions. It is worth noting that individual unfairness can be evaluated at the level of each example, whereas group fairness metrics are defined over sets of validation or test examples. To compute the fairness metrics, we randomly sample a subset of examples for each bias type such that each social group contributes an equal number of examples. Further details on test set sampling are provided in Appendix A.
@@
We employ 14 variants of commonly used input-based post-hoc explanation methods, covering diverse methodological categories: Attention (Bahdanau et al., 2015), Attention rollout (Attn rollout, Abnar and Zuidema, 2020), Attention flow (Attn flow, Abnar and Zuidema, 2020), Gradient (Grad,  Simonyan et al., 2014), Input x Gradient (IxG,  Kindermans et al., 2016), Integrated Gradients (IntGrad,  Sundararajan et al., 2017), Occlusion (Li et al., 2016), DeepLift (Shrikumar et al., 2017), and KernelSHAP (Lundberg and Lee, 2017). For methods that attribute predictions to embeddings, we aggregate attribution scores into a single feature importance value using either the mean or the L2 norm. For Occlusion, we additionally report results obtained by taking the absolute value of each attribution score prior to computing sensitive token reliance scores (denoted as Occlusion abs). We also study rationales generated by LLMs and find that these rationales are not as reliable as input-based explanations in detecting bias (Section 6).
@@
Qwen3 Zero-shot 69.55/79.75/77.50 0.60/0.00/17.40 7.13/1.40/21.07 13.25/3.71/5.17 2.55/2.41/3.32
Few-shot 63.00/76.40/74.30 1.90/1.60/20.80 10.17/4.77/26.67 8.04/7.21/8.22 3.30/4.03/4.60
@@ 4 Quantitative Analyses of Fairness and Explainability
To comprehensively understand the relationship between explainability and fairness in NLP models, we examine three ways in which model explanations can be applied to promote fairness. The subsequent sections detail the experimental setups for each application and report the corresponding results. For brevity, we report results on Civil Comments using BERT trained on single bias types and Qwen3. Results for additional models and the Jigsaw dataset are presented in Appendix B to E.
@@ 4.1 Model Performance and Fairness
As a prerequisite, we first summarize the performance and fairness of the evaluated models. The results in Table 1 show that no single debiasing method consistently improves all fairness metrics. For BERT and Qwen3, CDA and fairness instruction achieve the largest reductions in individual unfairness, yet they may simultaneously amplify biases on other metrics. Other debiasing methods show a similar pattern: they reduce bias for a specific metric or bias type, but the improvement does not generalize across different setups. These limitations underscore the importance of leveraging explanations for bias detection and mitigation. We find similar results for other models and for Jigsaw, which we provide in Appendix B along with a discussion on model performance and fairness.
@@ 4.2 RQ1: Explanations For Bias Detection
We present results for default and debiased models where individual unfairness remains high after debiasing, as bias detection is particularly critical in these cases. Specifically, we report results for models with the highest average $\text{Avg}_{\text{iu}}$ scores across bias types, namely default, group balance, and causal debias for BERT, and zero-shot, few-shot, and fairness imagination prompting for Qwen3. Results for religion as well as other models and the Jigsaw dataset are provided in Appendix C.
@@ Results
Figure 1 shows that the best-performing explanation methods, such as Grad L2, IxG L2, DeepLift L2, Occlusion, and Occlusion abs, generally achieve high fairness correlations across different models and bias types, indicating a strong ability to detect biased predictions. Besides, their fairness correlations are mostly statistically significant ($p<\alpha=0.05)$ in all, or in all but one, class-group categories, which confirms their reliability. Among these methods, Occlusion and Occlusion abs perform best with BERT models, whereas the L2-based methods Grad L2, IxG L2, and DeepLift L2 are most effective with Qwen3.
When comparing different variants of the same explanation family, mean-based approaches perform considerably worse than their L2-based counterparts, and also underperform compared to undirected attention-based methods. We attribute this limitation to their dependence on accurately determining the direction of each token’s contribution, a challenge that attention- and L2-based explanations do not face. Our analysis in Appendix G further shows that the effectiveness of explanation-based bias detection is not determined by explanation faithfulness, underscoring the need for careful evaluation when selecting methods for bias identification.
@@ 4.3 RQ2: Explanations for Model Selection
Our experiments consist of three steps: (1) for all default and debiased models (seven encoder-only and four decoder-only), we generate predictions on a validation set and compute explanation-based metrics; (2) we compute fairness metrics on the test set for each model; and (3) we evaluate model selection ability using two measures: Spearman’s rank correlation ($\rho$) between validation set explanation-based metrics and test set fairness metrics, which reflects the ability to rank models, and mean reciprocal rank of the fairest model (MRR@1), which reflects the ability to select the fairest model. Higher correlations and MRR@1 indicate that an explanation method is useful for ranking models and selecting the fairest one. Specifically, we use the average absolute sensitive token reliance on the validation set as the explanation-based metric to rank and select models based on average individual unfairness on the test set. [^2] We have evaluated other metrics to predict group fairness outcomes. However, neither explanation-based metrics nor validation set fairness achieved rank correlations beyond random chance with the test set results. As a baseline, we report results of using the validation set average individual unfairness as the predictor of test set fairness performance. The results are averaged over six and three random validation set selections for encoder- and decoder-only models, respectively. Results for more models and the Jigsaw dataset are presented in Appendix D.

### v2

@@ Abstract
Natural language processing (NLP) models often replicate or amplify social bias from training data, raising concerns about fairness. At the same time, their black-box nature makes it difficult for users to recognize biased predictions and for developers to effectively mitigate them. While some studies suggest that input-based explanations can help detect and mitigate bias, others question their reliability in ensuring fairness. Existing research on explainability in fair NLP has been predominantly qualitative, with limited large-scale quantitative analysis. In this work, we conduct the first systematic study of the relationship between explainability and fairness in hate speech detection, focusing on both encoder- and decoder-only models. We examine three key dimensions: (1) identifying biased predictions, (2) selecting fair models, and (3) mitigating bias during model training. Our findings show that input-based explanations can effectively detect biased predictions and serve as useful supervision for reducing bias during training, but they are unreliable for selecting fair models among candidates. Our code is available at https://github.com/Ewanwong/fairness_x_explainability.
@@ 1 Introduction
Meanwhile, current NLP models are mostly based on black-box neural networks. Despite their strong capacities, the complex architecture and large number of parameters of these models make it hard for humans to understand their behaviors (Bommasani et al., 2021). To understand neural NLP models, different types of explanations have been devised, such as input-based explanations (Yin and Neubig, 2022; Deiseroth et al., 2023; Madsen et al., 2024; Wang et al., 2025b), natural language explanations (Ramnath et al., 2024; Wang et al., 2025a), and concept-based explanations (Yu et al., 2024; Raman et al., 2024). Among these, input-based explanations, often referred to as rationales, indicate the contribution of each token to models’ predictions, and thus provide the most direct insights into models’ behaviors (Arras et al., 2019; Atanasova et al., 2022; Lyu et al., 2024).
@@
Our experiments demonstrate that input-based explanations can effectively detect biased predictions (RQ1), are less reliable for automatic fair model selection (RQ2), and can help reduce bias during model training (RQ3). Furthermore, our analyses indicate that explanation-based bias detection remains robust even when models are trained to reduce reliance on sensitive features, and that these explanations outperform LLM judgments in identifying bias.
@@ 2 Related Work
Input-based explanations in NLP models aim to attribute model predictions to each input token (Lyu et al., 2024). They can be broadly categorized based on how they generate explanations: gradient-based (Simonyan et al., 2014; Kindermans et al., 2016; Sundararajan et al., 2017; Enguehard, 2023), propagation-based (Bach et al., 2015; Shrikumar et al., 2017; Ferrando et al., 2022; Modarressi et al., 2022; 2023), perturbation-based (Li et al., 2016; Ribeiro et al., 2016; Lundberg and Lee, 2017; Deiseroth et al., 2023), and attention-based methods (Bahdanau et al., 2015; Abnar and Zuidema, 2020). While most prior work has focused on encoder-only models, recent studies have also explored explaining the behaviors of generative models (Yin and Neubig, 2022; Ferrando et al., 2022; Enouen et al., 2024; Cohen-Wang et al., 2024).
@@
Another line of research focuses on mitigating bias with explanations (Dimanov et al., 2020; Kennedy et al., 2020; Rao et al., 2023; Liu et al., 2024). For example, Hickey et al. (2020) improves fairness by reducing reliance on sensitive features during training with SHAP explanations. Bhargava et al. (2020) and González-Silot et al. (2025) first identify predictive sensitive features using LIME and SHAP, respectively, and then remove them prior to model training. Relatedly, Grabowicz et al. (2022) traces unfairness metrics to input features and adjusts them to mitigate bias.
However, recent research has challenged the assumption that input-based explanations can be reliably used to detect and mitigate bias. First, current explanation methods may be unfaithful, meaning that they may not always reflect the true decision-making process of models (Jain and Wallace, 2019; Ye et al., 2025). This makes it difficult to reliably detect the use of sensitive features in predictions. Second, efforts to reduce the influence of sensitive features can lead to unintended consequences, sometimes degrading both task performance and fairness of models (Dimanov et al., 2020). Finally, models can be deliberately trained to assign lower importance to sensitive features, thereby masking biased predictions when explanations are inspected (Slack et al., 2020; Pruthi et al., 2020).
Despite growing interest in this topic, most existing work remains qualitative or restricted to limited setups (Balkir et al., 2022; Deck et al., 2024). To the best of our knowledge, this is the first study to quantitatively and comprehensively examine the relationship between explainability and fairness in NLP models. We focus on hate speech detection as a particularly critical application. Prior research has shown that biased NLP models often rely on demographic information such as race and gender, leading to inferior performance on marginalized groups in this task (Sap et al., 2019; Mathew et al., 2021). Detecting and mitigating such biased behaviors are therefore essential for equitable social media participation across all groups. We define hate speech and social bias and review related fairness and explainability work in hate speech detection in Appendix A.
Figure 1: Workflow diagram illustrating the processes used to address each research question. Sensitive tokens are shown in blue boxes, and the intensity of the green shading reflects each word’s contribution to the model’s prediction.
@@ 3 Experimental Setup
An input-based explanation assigns an attribution score to each token in $\mathbf{x}$ for class $c$: $a_{1}^{c},a_{2}^{c},\dots,a_{n}^{c}$}, indicating their contribution to the prediction of class $c$. Following Dimanov et al. (2020), we compute attribution scores on the sensitive tokens, $a_{g_{1}}^{c},a_{g_{2}}^{c},\dots,a_{g_{m}}^{c}$}, which we refer to as the sensitive token reliance scores. To handle cases where multiple sensitive tokens appear in the same sentence, we take the maximum absolute attribution value as the reliance score for that example [^1] We have experimented with normalizing feature importance scores but found that using raw scores yielded the best results. We also evaluated sum and average aggregation methods beyond taking the max absolute value and observed similar outcomes.:
@@
We use two hate speech detection datasets: Civil Comments (Borkan et al., 2019) and Jigsaw (cjadams et al., 2019). To ensure coverage, we focus on three bias types and their associated groups: race (black/white), gender (female/male), and religion (Christian/Muslim/Jewish). We include examples containing identity-marking terms but exclude those with derogatory or slur-based references, as the latter can reasonably serve as direct evidence for toxic predictions. The sensitive token vocabulary is derived from Caliskan et al. (2017) and Wang and Demberg (2024a). Further details on dataset pre-processing are provided in Appendix B.
@@
We evaluate two major classes of NLP models: encoder-only models (BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)) and decoder-only large language models (Llama3.2-3B-Instruct (Dubey et al., 2024), Qwen3-4B, and Qwen3-8B (Yang et al., 2025a), all of which are aligned to human values). We fine-tune encoder-only models on data subsets that either target a single bias type or combine all bias types. For decoder-only models, we use an instruction-based setup where the model is prompted to decide whether a test example contains hate speech. The prompt includes the definition of hate speech, the test example, and a corresponding question. As a baseline, we adopt the zero-shot setting as the default configuration.
Beyond conventional fine-tuning and prompting, we also investigate the interaction between explainability and fairness in debiased models. For encoder-only models, we apply pre-processing techniques such as group balance (Kamiran and Calders, 2012), group-class balance (Dixon et al., 2018), and counterfactual data augmentation (CDA, Zmigrod et al., 2019), as well as in-processing techniques including dropout (Webster et al., 2020), attention entropy (Attanasio et al., 2022), and causal debias (Zhou et al., 2023). For decoder-only models, we incorporate bias reduction through prompt design, including few-shot, fairness imagination (Chen et al., 2025), and fairness instruction prompting (Chen et al., 2025). We do not include reasoning models and chain-of-thought prompting, as we find that their predictions are primarily attributed to intermediate reasoning steps rather than the input text, which complicates analysis and falls beyond the scope of this work. Further details are provided in Appendix B.
@@
For both types of metrics, higher scores indicate more bias in model predictions. It is worth noting that individual unfairness can be evaluated at the level of each example, whereas group fairness metrics are defined over sets of validation or test examples. To compute the fairness metrics, we randomly sample a subset of examples for each bias type such that each social group contributes an equal number of examples. Further details on test set sampling are provided in Appendix B.
@@
We employ 16 variants of commonly used input-based post-hoc explanation methods, selected to represent a diverse range of methodological categories: Attention (Bahdanau et al., 2015), Attention rollout (Attn rollout, Abnar and Zuidema, 2020), Attention flow (Attn flow, Abnar and Zuidema, 2020), Gradient (Grad,  Simonyan et al., 2014), Input x Gradient (IxG, Kindermans et al., 2016), Integrated Gradients (IntGrad,  Sundararajan et al., 2017), Occlusion (Li et al., 2016), DeepLift (Shrikumar et al., 2017), KernelSHAP (Lundberg and Lee, 2017), DecompX (Modarressi et al., 2023), and Progressive Inference (ProgInfer, Kariyappa et al., 2024) [^2] We apply DecompX only to encoder-only models and Progressive Inference only to decoder-only models, following the setups of the original papers.. For methods that attribute predictions to embeddings, we aggregate attribution scores into a single feature importance value using either the mean or the L2 norm. For Occlusion, we additionally report results obtained by taking the absolute value of each attribution score prior to computing sensitive token reliance scores (denoted as Occlusion abs). The time and GPU memory costs for each method are shown in Appendix C. We also study rationales generated by LLMs and find that they are not as reliable as input-based explanations in detecting bias (Section 6).
@@
Qwen3-4B Zero-shot 69.55/79.75/77.50 0.60/0.00/17.40 7.13/1.40/21.07 13.25/3.71/5.17 2.55/2.41/3.32
Few-shot 70.15/80.73/79.53 1.80/0.65/18.93 10.02/2.50/19.31 11.89/9.15/5.57 3.18/3.34/3.76
@@ 4 Quantitative Analyses of Fairness and Explainability
To comprehensively understand the relationship between explainability and fairness in NLP models, we examine three ways in which model explanations can be applied to promote fairness. The subsequent sections detail the experimental setups for each application and report the corresponding results. The workflow for our research questions is shown in Figure 1. For brevity, we report results on Civil Comments using BERT trained on single bias types and Qwen3-4B. We further demonstrate that our findings extend to other models and the Jigsaw dataset (Appendix DG), and that they generalize across tasks, model alignment types, and sensitive token sets beyond those considered in the main experiments (Appendix H).
@@ 4.1 Model Performance and Fairness
As a prerequisite, we first summarize the performance and fairness of the evaluated models. The results in Table 1 show that no single debiasing method consistently improves all fairness metrics. For BERT and Qwen3-4B, CDA and fairness instruction achieve the largest reductions in individual unfairness, yet they may simultaneously amplify biases on other metrics. Other debiasing methods show a similar pattern: they reduce bias for a specific metric or bias type, but the improvement does not generalize across different setups. These limitations underscore the importance of leveraging explanations for bias detection and mitigation. We find similar results for other models and for Jigsaw, which we provide in Appendix D along with a discussion on model performance and fairness.
@@ 4.2 RQ1: Explanations For Bias Detection
We present results for default and debiased models where individual unfairness remains high after debiasing, as bias detection is particularly critical in these cases. Specifically, we report results for models with the highest average $\text{Avg}_{\text{iu}}$ scores across bias types, namely default, group balance, and causal debias for BERT, and zero-shot, few-shot, and fairness imagination prompting for Qwen3-4B. Results for religion as well as other models and the Jigsaw dataset are provided in Appendix E.
@@ Results
Figure 2 shows that the best-performing explanation methods, such as Grad L2, IxG L2, DeepLift L2, Occlusion, and Occlusion abs, generally achieve high fairness correlations across different models and bias types, indicating a strong ability to detect biased predictions. Besides, their fairness correlations are mostly statistically significant ($p<\alpha=0.05)$ in all, or in all but one, class-group categories, which confirms their reliability. Among these methods, Occlusion and Occlusion abs perform best with BERT models, whereas the L2-based methods Grad L2, IxG L2, and DeepLift L2 are most effective with Qwen3-4B.
When comparing different variants of the same explanation family, mean-based approaches perform considerably worse than their L2-based counterparts, and also underperform compared to undirected attention-based methods. We attribute this limitation to their dependence on accurately determining the direction of each token’s contribution, a challenge that attention- and L2-based explanations do not face. Our analysis in Appendix I further shows that the effectiveness of explanation-based bias detection is not determined by explanation faithfulness, underscoring the need for careful evaluation when selecting methods for bias identification. In addition, our human study in Appendix J shows that explanation-based signals can assist human fairness auditing.
@@ 4.3 RQ2: Explanations for Model Selection
Our experiments consist of three steps: (1) for all default and debiased models (seven encoder-only and four decoder-only), we generate predictions on a validation set and compute explanation-based metrics; (2) we compute fairness metrics on the test set for each model; and (3) we evaluate model selection ability using two measures: Spearman’s rank correlation ($\rho$) between validation set explanation-based metrics and test set fairness metrics, which reflects the ability to rank models, and mean reciprocal rank of the fairest model (MRR@1), which reflects the ability to select the fairest model. Higher rank correlations and MRR@1 indicate that an explanation method is useful for ranking models and selecting the fairest one. Specifically, we use the average absolute sensitive token reliance on the validation set as the explanation-based metric to rank and select models based on average individual unfairness on the test set. [^3] We have evaluated other metrics to predict group fairness outcomes. However, neither explanation-based metrics nor validation set fairness achieved rank correlations beyond random chance with the test set results. The full set of evaluated metrics is provided in Appendix F.